Nomic InitialsetofrulesofNomic Immutablerules 101Allplayersmustalwaysabidebyalltherulesthenineffect,intheforminwhichtheyare thenineffect.TherulesintheInitialSetareineffectwheneveragamebegins.TheInitial SetconsistsofRules101-116(immutable)and201-213(mutable). 102Initiallyrulesinthe100’sareimmutableandrulesinthe200’saremutable.Rules subsequentlyenactedortransmuted(thatis,changedfromimmutabletomutableorvice versa)maybeimmutableormutableregardlessoftheirnumbers,andrulesintheInitialSet maybetransmutedregardlessoftheirnumbers. 103Arule-changeisanyofthefollowing: 1.theenactment,repeal,oramendmentofamutablerule; 2.theenactment,repeal,oramendmentofanamendmentofamutablerule;or 3.thetransmutationofanimmutableruleintoamutableruleorviceversa. (Note:Thisdefinitionimpliesthat,atleastinitially,allnewrulesaremutable;immutable rules,aslongastheyareimmutable,maynotbeamendedorrepealed;mutablerules,as longastheyaremutable,maybeamendedorrepealed;anyruleofanystatusmaybe transmuted;noruleisabsolutelyimmunetochange.) 104Allrule-changesproposedintheproperwayshallbevotedon.Theywillbeadoptedifand onlyiftheyreceivetherequirednumberofvotes. 105Everyplayerisaneligiblevoter.Everyeligiblevotermustparticipateineveryvoteonrulechanges. 106Allproposedrule-changesshallbewrittendownbeforetheyarevotedon.Iftheyare adopted,theyshallguideplayintheforminwhichtheywerevotedon. 107Norule-changemaytakeeffectearlierthanthemomentofthecompletionofthevotethat adoptedit,evenifitswordingexplicitlystatesotherwise.Norule-changemayhave retroactiveapplication. 108Eachproposedrule-changeshallbegivenanumberforreference.Thenumbersshallbegin with301,andeachrule-changeproposedintheproperwayshallreceivethenextsuccessive integer,whetherornottheproposalisadopted. Ifaruleisrepealedandreenacted,itreceivesthenumberoftheproposaltoreenactit.Ifa ruleisamendedortransmuted,itreceivesthenumberoftheproposaltoamendor transmuteit.Ifanamendmentisamendedorrepealed,theentireruleofwhichitisapart receivesthenumberoftheproposaltoamendorrepealtheamendment. 109Rule-changesthattransmuteimmutablerulesintomutablerulesmaybeadoptedifandonly ifthevoteisunanimousamongtheeligiblevoters.Transmutationshallnotbeimplied,but mustbestatedexplicitlyinaproposaltotakeeffect. 110Inaconflictbetweenamutableandanimmutablerule,theimmutableruletakesprecedence andthemutableruleshallbeentirelyvoid.Forthepurposesofthisruleaproposalto transmuteanimmutableruledoesnot“conflict”withthatimmutablerule. 111Ifarule-changeasproposedisunclear,ambiguous,paradoxical,ordestructiveofplay,orif itarguablyconsistsoftwoormorerule-changescompoundedorisanamendmentthat makesnodifference,orifitisotherwiseofquestionablevalue,thentheotherplayersmay suggestamendmentsorargueagainsttheproposalbeforethevote.Areasonabletimemust beallowedforthisdebate.Theproponentdecidesthefinalforminwhichtheproposalisto bevotedonand,unlesstheJudgehasbeenaskedtodoso,alsodecidesthetimetoend debateandvote. 1 Nomic 112Thestateofaffairsthatconstituteswinningmaynotbealteredfromachievingnpointsto anyotherstateofaffairs.Themagnitudeofnandthemeansofearningpointsmaybe changed,andrulesthatestablishawinnerwhenplaycannotcontinuemaybeenactedand (whiletheyaremutable)beamendedorrepealed. 113Aplayeralwayshastheoptiontoforfeitthegameratherthancontinuetoplayorincura gamepenalty.Nopenaltyworsethanlosing,inthejudgmentoftheplayertoincurit,may beimposed. 114Theremustalwaysbeatleastonemutablerule.Theadoptionofrule-changesmustnever becomecompletelyimpermissible. 115Rule-changesthataffectrulesneededtoalloworapplyrule-changesareaspermissibleas otherrule-changes.Evenrule-changesthatamendorrepealtheirownauthorityare permissible.Norule-changeortypeofmoveisimpermissiblesolelyonaccountoftheself-referenceorself-applicationofarule. 116Whateverisnotprohibitedorregulatedbyaruleispermittedandunregulated,withthe soleexceptionofchangingtherules,whichispermittedonlywhenaruleorsetofrules explicitlyorimplicitlypermitsit. Mutablerules 201Playersshallalternateinclockwiseorder,takingonewholeturnapiece.Turnsmaynotbe skippedorpassed,andpartsofturnsmaynotbeomitted.Allplayersbeginwithzero points. Inmailandcomputergames,playersshallalternateinalphabeticalorderbysurname. 202Oneturnconsistsoftwopartsinthisorder: 1.proposingonerule-changeandhavingitvotedon,and 2.throwingonedieonceandaddingthenumberofpointsonitsfacetoone’sscore. Inmailandcomputergames,insteadofthrowingadie,playerssubtract291fromthe ordinalnumberoftheirproposalandmultiplytheresultbythefractionoffavorablevotesit received,roundedtothenearestinteger.(Thisyieldsanumberbetween0and10forthe firstplayer,withtheupperlimitincreasingbyoneeachturn;morepointsareawardedfor morepopularproposals.) 203Arule-changeisadoptedifandonlyifthevoteisunanimousamongtheeligiblevoters.If thisruleisnotamendedbytheendofthesecondcompletecircuitofturns,itautomatically changestorequireonlyasimplemajority. 204Ifandwhenrule-changescanbeadoptedwithoutunanimity,theplayerswhovoteagainst winningproposalsshallreceive10pointseach. 205Anadoptedrule-changetakesfulleffectatthemomentofthecompletionofthevotethat adoptedit. 206Whenaproposedrule-changeisdefeated,theplayerwhoproposeditloses10points. 207Eachplayeralwayshasexactlyonevote. 208Thewinneristhefirstplayertoachieve100(positive)points. Inmailandcomputergames,thewinneristhefirstplayertoachieve200(positive)points. 209Atnotimemaytherebemorethan25mutablerules. 210Playersmaynotconspireorconsultonthemakingoffuturerule-changesunlesstheyare team-mates. Thefirstparagraphofthisruledoesnotapplytogamesbymailorcomputer. 2 Nomic 211Iftwoormoremutablerulesconflictwithoneanother,oriftwoormoreimmutablerules conflictwithoneanother,thentherulewiththelowestordinalnumbertakesprecedence. Ifatleastoneoftherulesinconflictexplicitlysaysofitselfthatitdeferstoanotherrule(or typeofrule)ortakesprecedenceoveranotherrule(ortypeofrule),thensuchprovisions shallsupersedethenumericalmethodfordeterminingprecedence. Iftwoormorerulesclaimtotakeprecedenceoveroneanotherortodefertooneanother, thenthenumericalmethodagaingoverns. 212Ifplayersdisagreeaboutthelegalityofamoveortheinterpretationorapplicationofarule, thentheplayerprecedingtheonemovingistobetheJudgeanddecidethequestion. Disagreementforthepurposesofthisrulemaybecreatedbytheinsistenceofanyplayer. ThisprocessiscalledinvokingJudgment. WhenJudgmenthasbeeninvoked,thenextplayermaynotbeginhisorherturnwithout theconsentofamajorityoftheotherplayers. TheJudge’sJudgmentmaybeoverruledonlybyaunanimousvoteoftheotherplayers takenbeforethenextturnisbegun.IfaJudge’sJudgmentisoverruled,thentheplayer precedingtheJudgeintheplayingorderbecomesthenewJudgeforthequestion,andsoon, exceptthatnoplayeristobeJudgeduringhisorherownturnorduringtheturnofa team-mate. UnlessaJudgeisoverruled,oneJudgesettlesallquestionsarisingfromthegameuntilthe nextturnisbegun,includingquestionsastohisorherownlegitimacyandjurisdictionas Judge. NewJudgesarenotboundbythedecisionsofoldJudges.NewJudgesmay,however,settle onlythosequestionsonwhichtheplayerscurrentlydisagreeandthataffectthecompletion oftheturninwhichJudgmentwasinvoked.AlldecisionsbyJudgesshallbeinaccordance withalltherulesthenineffect;butwhentherulesaresilent,inconsistent,orunclearonthe pointatissue,thentheJudgeshallconsidergame-customandthespiritofthegamebefore applyingotherstandards. 213Iftherulesarechangedsothatfurtherplayisimpossible,orifthelegalityofamovecannot bedeterminedwithfinality,orifbytheJudge’sbestreasoning,notoverruled,amove appearsequallylegalandillegal,thenthefirstplayerunabletocompleteaturnisthe winner. Thisruletakesprecedenceovereveryotherruledeterminingthewinner. 3